Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page - 20130221

Observers: Nancy Elias-Rosa, Antonia Morales-Garoffolo
Data reduction team: Sandra Benitez, Stefan Taubenberger

Observations with SOFI.  
**Humidity from 25% to 4% (!!!!) to the end of the night.

Calibration data done in the early afternoon.

**We have just had an electricity cut (UT 20:25). NTT and SOFI seem to be responding correctly. 

UT 00:28      SN2012ec                      airmass=1.52        JHKs                 

--DIMM SEEING 0.65"--

UT 00:55      Hip036187                     airmass=1.62        BG,RG                Telluric STD for OGLE-2012-SN-006.

UT 01:10      OGLE-2012-SN-006              airmass=1.57        BG,RG                Observation aborted (UT 01:43), not any trace of the SN in the first AB cycle in BG. We start again (UT 01:49).  

UT 04:30      Hip014112                     airmass=2.27        BG,RG                Telluric STD for OGLE-2012-SN-006.

--DIMM SEEING 0.9"--

UT 04:39      OGLE-2012-SN-006              airmass=2.28        RG                   One more AB cycle.

UT 05:08      OGLE-2012-SN-006              airmass=2.44        JHKs

--DIMM SEEING 0.79"--

UT 05:54      SN2013K                       airmass=2.07        JHKs

UT 06:44      GD153                         airmass=1.64        BG,RG                Flux STD.

--DIMM SEEING 0.69"--

UT 06:58      Hip045705                     airmass=1.28        BG                   Telluric STD for SN2012hs.            

--DIMM SEEING 0.59"--

UT 07:10      SN2012hs                      airmass=1.26        BG

UT 08:11      PSNJ15213775-0722183          airmass=1.19        JHKs

UT 09:30      END of the night


Comments